# Stagewise Reinforcement Learning and the Geometry of the Regret Landscape

### v1

@@ 2.6 The Local Learning Coefficient
where $U$ is a sufficiently small open neighborhood of $w_{0}$. This formula describes the concentration of the posterior in a region $U$ in terms of two invariants: the empirical regret $G_{n}(w_{0})$ (lower regret means lower $F_{n}(U)$, hence higher $Z_{n}(U)$ and thus posterior concentration) and the learning coefficient $\lambda$ (lower $\lambda$ means a simpler parametrized policy, and higher posterior concentration). The prior $\phi(w)$ contributes to the free energy, and thus posterior concentration, only through constant order terms. These can certainly play a significant role for low $n$, but we do not consider this explicitly there.
2.7 Implications of the free energy formula
The central insight for Bayesian learning in deep RL that we draw from the free energy formula (8) is that learning consists of an evolving tradeoff between regret and complexity. This is the translation to the setting of RL of a perspective due to Watanabe (2009, §7.6), see also (Balasubramanian, 1997; Chen et al., 2024b).
@@ 2.7 Implications of the free energy formula
Note that since the relationship between the free energy and the posterior is exponential, these switches in the sign of the free energy as a function of $n$ can translate to rapid changes in posterior concentration. Following Watanabe (2018) and Chen et al. (2024b) we refer to these sudden shifts in the concentration of the posterior as Bayesian phase transitions.
This suggests a picture of the Bayesian learning process as a series of phase transitions, where the posterior jumps between neighborhoods of local minima $w_{i}$ of $G$, according to a changing tradeoff between $G$ and $\lambda$ in the free energy formula. The most typical transition decreases $G$ and increases $\lambda$ (the posterior is willing, at some dataset size, to pay the free-energetic cost of higher complexity to get lower regret).
This is not that surprising if we think of learning as the process of incorporating information from observed trajectories into the network parameter, but framed differently it has a counter-intuitive implication, flagged already in the introduction: given any finite number of observed trajectories, a more optimal policy (lower regret) may not be more optimal from a Bayesian perspective (lower free energy) according to the concentration of the generalized posterior. This has implications for AI alignment, which we review in Section 5.3.
@@ 3.1 Summary
During training we see stagewise development in policy space: the model transitions between plateaus in which the regret is roughly constant, the policy remains in a fixed region with qualitatively consistent behavior and the transitions between plateaus are relatively rapid. We refer to periods where the regret is approximately constant as phases and transitions between them as phase transitions (and one might compare this phenomenon to the phase transition seen in a toy model of supervised learning in Chen et al. (2024b)). For example, we will discuss some of the following common phases in policy space below.
@@
    We see gradual decrease in the LLC estimate within phases 2 and 3 – the policy, and therefore the regret, does not change significantly but the representation of the policy within parameter space does change. The LLC of phase 1 exhibits a consistent value of $41.5\pm 8.2$ during training and across values of $\alpha$ and $\gamma$.
  3. 3.
    The LLC detects phase transitions even when they are not visible from the loss alone. Indeed, we may estimate the LLC in states corresponding the the $\alpha=0$ distribution in which the goal is always in the top-left corner. Even if only these states are seen during estimation we still see rapid increases in the LLC estimator upon the phase transition.
@@ 3.2.2 Model
The architecture of the model directly follows that used for Abdel Sadek et al. (2025), which itself is a modified version of the Large architecture described in Espeholt et al. (2018): A convolutional network with 15 convolutional layers, but the LSTM that normally follows is replaced by a simple feedforward network that takes both the output from the convolutional block, and a one-hot encoding of the previous action. Full details of the architecture are given in Espeholt et al. (2018).
Note that the inclusion of the previous action makes the model Markovian as a function of a state and previous action pair, $a_{t}\sim\pi(\cdot|s_{t},a_{t-1})$.
@@
Given the simplicity of the environment and given that there are only $121\times 120\times 5=72600$ possible combinations [^4] At the start of an episode, there is no previous action, so $a_{t-1}$ can take on 5 different values of state and previous action $(s_{t},a_{t-1})$ we can analytically compute the optimal reward $R_{\mathrm{max}}$ and the exact reward $R(w)$ of a given current policy $\pi_{w}$. This allows us to analytically compute the regret, $G(w)=R_{\mathrm{max}}-R(w)$ during various stages of training, as well as visualize the policy for each possible location of the agent given a goal location.
@@ 3.3 Details of LLC Estimation
We define an estimator for the local learning coefficient (see Definition 2.6) using the preconditioned stochastic gradient Langevin dynamics (pSGLD) algorithm Li et al. (2015) to estimate sampling from the Gibbs posterior. This combines RMSNorm-style adaptive step sizes with SGLD Welling and Teh (2011). Rather than imposing the sharp cutoff associated to the restriction to a neighborhood we use a Gaussian prior centered at $w^{*}$ with variance $\sigma^{2}$. The Gibbs posterior then takes the form
@@
The hyperparameters are the sample size $n$, the inverse temperature $\beta$, which controls the contribution of the regret, and the variance $\sigma^{2}$, which controls proximity to $w^{*}$}. For more details on these hyperparameters, see Watanabe (2013); Lau et al. (2025); Hoogland et al. (2025). In loc. cit. the inverse of the variance $\sigma^{2}$ is referred to as the localization strength. Details of the hyperparameters used for LLC estimation can be found in Appendix B.
@@ 3.4 Phase Transitions
A training run with $\alpha=0.68,\gamma=0.99$ is shown in Figure 1. We plot the regret, the difference between the value of the current policy, and that of an optimal policy, as a function of the number of environment steps seen during training. Three chosen points during training are illustrated with depictions of the learned policy in each phase.
@@ 3.4.1 Automatic Phase Detection
The value of $\delta$ is chosen to be smaller than the normalized distance between phase 1 and phase 2b in order to detect the transition between those phases. The precise choice of $\delta=0.15$ is not critical for the location of the phase transitions, as there is a range over which we could choose $\delta$ with only minor changes to the corresponding locations of phase transitions (see Appendix I).
@@ 3.4.2 Distribution of Phase Transitions
The training trajectory – and therefore the time steps at which phase transitions occur – are random variables, in this section we discuss their empirical distribution. Specifically we will present statistics in eight cases: where the model was trained with $\alpha=0.47$ or $\alpha=0.68$ and with $\gamma=0.9,0.95.0.975$ and $0.99$. In each hyperparameter case we trained models with these hyperparameters and $\sim 25$ independent random seeds. In Figure 3 we exhibit the distribution of the training steps at which phase transitions occur as a function of $\gamma$ for each of the two sampled $\alpha$-values. The models transition into phase 1 consistently early in training, the transition from phase 1 to phase 2b depends on $\gamma$ as we will discuss below.
@@
    The number of training steps to transition into $\pi^{1}$ is very fast, and unrelated to the choice of $\alpha$$\gamma$ presented.
@@
    The higher the discount rate, the longer it takes for the model to converge to $\pi^{2b}$, if at all. For $\alpha,\gamma=0.68,0.99$, the policy $\pi^{2b}$ is never observed, but is skipped entirely, transitioning from $\pi^{1}\to\pi^{3}$ directly without visiting $\pi^{2b}$ first. For $\alpha,\gamma=0.68,0.99$, during training the policy would sometimes, but not always, skip $\pi^{2b}$ and transition directly from $\pi^{1}\to\pi^{3}$. This is noted with the * to indicate we are not measuring the same quantity as the other runs.
  3. 3.
    For $\alpha,\gamma=0.47,0.99$, some of the runs never reached $\pi^{3}$ by the time 5 billion environmental steps had been seen. This means that the collected statistics for this run should be considered with a grain of salt, as early termination of runs that may have converged later are not included in the plot, and this causes a downwards bias in the result.
We can also try to predict the phase transitions based on the hyperparameters with which it was trained, see Appendix C.
@@ 3.4.3 LLC Estimates and Phases
Figure 1 shows the LLC estimate and regret across training for a representative model trained with $\alpha=0.68$ and $\gamma=0.975$, exhibiting two phase transitions. We present all the estimated LLC curves in Appendix G, Figure 10. We see rapid increases in the LLC estimate at each phase transition, and the increasingly complex phases are associated with increasingly large LLC estimates. We also note a gradual decrease of the LLC estimates in the second halves of phases 2 and 3 as stochastic gradient descent converges to simpler representations of the corresponding local minimum.
@@
Next we discuss the distribution of the LLC estimates within each phase. We performed LLC estimation for 23 independent models trained with $\alpha=0.68$ and $\gamma=0.975$ – those hyperparameters being selected to maximize the proportion of runs exhibiting two phase transitions. We estimated the LLC at two checkpoints in each of the three phases: Two points at the 1/3 and 2/3 tertiles with respect to the logarithm of the checkpoint number. In Figure 4 we show the distribution of the LLC estimate at these 6 points during training. Using the two internal points within each phase we see mean LLC estimates of 30.92 in phase one (with sample variance $\hat{\sigma}=3.38$), 106.80 in phase two ($\hat{\sigma}=16.19$) and 561.52 in phase three ($\hat{\sigma}=40.49$).
@@
Figure 4: Distribution of average LLC estimates across phases. The LLC has been estimated on distribution with $\alpha=0.68$ and $\gamma=0.975$.
3.4.4 LLC Estimates with $\alpha=0$
Let us now examine LLC estimates for the same models as those described in the section above, but estimated with fixed values of $\alpha$ and $\gamma$ (as described in Remark 3.2 above). We in particular discuss the behavior with $\alpha=0$. In other words during LLC estimation we only consider the behavior of the model in states where the goal is in the top-left corner. Note that in such states we cannot distinguish between phases 2 and 3 by means of the regret anymore: these phases are both optimal and receive zero regret. Nevertheless LLC estimation still allows these phases to be distinguished. We use a fixed value $\gamma=0.98$ in this section (and refer to Appendix G for LLC estimate data across models more broadly.)
We again show a representative LLC estimate performed with $\alpha=0$ in Figure 5. This estimation was again performed for a representative model trained with $\alpha=0.68$ and $\gamma=0.975$ exhibiting two phase transitions; the same model used in the previous section. Likewise in Figure 6 we again show the distribution of the LLC estimates at four checkpoints within each phase for 23 independent models trained with $\alpha=0.68$ and $\gamma=0.975$. We now see mean LLC estimates of 44.56 in phase one (with sample variance $\hat{\sigma}=5.76$), 102.54 in phase two ($\hat{\sigma}=34.07$) and 245.72 in phase three ($\hat{\sigma}=149.16$).
Figure 5: LLC and regret over training for a model trained with $\alpha=0.68$ and $\gamma=0.975$. The LLC has been computed with $\alpha=0$ and $\gamma=0.98$. We show two regret curves, the orange one with the same $\alpha$ and $\gamma$ as was used for training, and the green one with the same $\alpha$ and $\gamma$ as was used for LLC estimation.
Figure 6: Distribution of average LLC estimates across phases. The LLC has been estimated with $\alpha=0$ and $\gamma=0.98$.
4 Related Work
Measurements of complexity, including those quantifying flatness of a local minimum in the loss landscape have been widely studied for neural networks Hochreiter and Schmidhuber (1997); Keskar et al. (2016); Neyshabur et al. (2017); Dherin et al. (2022). Likewise many authors have explored the idea of simplicity bias in neural networks as an implicit regularization affecting generalization Shah et al. (2020); Teney et al. (2025); Pérez et al. (2018); Morwani et al. (2023). The notion of the loss-complexity tradeoffs and transitions between different solution classes is studied in Singh et al. (2023); Wurgaft et al. (2025); Munn and Wei (2025); Carroll et al. (2025).
The direct study of complexity bias in reinforcement learning specifically has been less explored. The geometry of the reward landscape was explored visually in many examples in Sullivan et al. (2022), which demonstrated the appearance in low-dimensional cross-sections of “cliffs” and “plateaus” in the landscape. The work Boucher et al. (2024) studied the inductive bias introduced by the technique of maximum entropy learning Ziebart et al. (2008) using a range of complexity measures including those that quantified local flatness of a minimum. This idea of flatness specifically as a bias favoring more robust optima was investigated in Lee and Yoon (2025). In Boopathy et al. (2023) the authors studied inductive biases in a range of contexts including reinforcement learning using an information-theoretic measure of complexity. Other authors have applied complexity bias to improve learning performance in reinforcement learning problems You et al. (2025); Lee et al. (2024), while others have studied the generalization of RL agents in grid world similar to the environments considered in this paper (Ramanauskas and Özgür Şimşek, 2023; Abdel Sadek et al., 2025).
We lastly mention some of the theoretical literature on the dynamics of reinforcement learning. The dynamics under policy gradient learning was studied in an illustrative model for reinforcement learning in Patel et al. (2025). The connection between complexity and stagewise development has recently been explored in the supervised learning case, particularly for large language models Wei et al. (2022a); Wang et al. (2024); Chen et al. (2024a). Finally in the neuroscience literature, it has been demonstrated that stagewise development also occurs during reinforcement learning in biology Liebana et al. (2025).
@@ 5 Discussion
5.1 Off-policy learning
$\bigl\{(w^{(k)},\tau^{(k)}_{j})\;:\;k\leq K,\;1\leq j\leq B_{k}\bigr\},$
i.e. it is a union of samples drawn under earlier iterates of the same SGD process. [^6] In distributed actor–learner architectures one should interpret $w^{(k)}$ as the parameters on the actor at the moment of rollout, which are typically a lagged copy of the learner parameters. Thus, even though our theory allows an arbitrary sequence $(w_{i})_{i=1}^{n}$, the “canonical” source of such a sequence in practice is simply the learning algorithm itself, via checkpoints or lagged actors.
which matches the on-policy training rule used in this paper (cf. (9)).
and replacing $g(\tau_{i})$ by its reward-to-go form yields the trajectory-sampled analogue of (15) (up to the overall sign corresponding to maximizing return vs. minimizing regret). Thus, stochastic gradient descent on $G_{n}$ corresponds to an off-policy REINFORCE-style update, while the generalized posterior is the Gibbs distribution associated to the same empirical objective.
5.2 Relation between Bayesian learning and SGD
In the theoretical part of this paper (Section 2) we have given a description of the Bayesian learning process in RL in terms of an evolving tradeoff between regret and complexity, as made precise by the free energy formula. This predicts Bayesian phase transitions as the generalized posterior $p(w|D_{n})\propto\exp(-nG_{n}(w))\phi(w)$ shifts its concentration between local minima of the population regret $G$. Here $n$ denotes the number of trajectories “seen” by the learner and incorporated into the posterior update.
In the experimental part of the paper (Section 3) we studied the training process of a deep RL agent by SGD. The optimization draws for each gradient step batches of $B$ trajectories and the gradient estimator in (9) is essentially $\nabla_{w}G_{n}(w)$ with $n=B$. After $S$ steps of gradient descent, information from $BS$ trajectories have influenced the current parameter.
To attempt to line these two learning processes up as much as possible, we can take the sequence of behavioral parameters of Section 5.1 to be $w_{1},\ldots,w_{1},\ldots,w_{S},\ldots,w_{S}$ where $w_{i}$ is the $i$th step of gradient descent and each parameter is repeated $B$ times, and take the $\tau_{i}$ to be the contents of mini-batches sampled during SGD. Thus in some sense $D_{n}$ is the set of trajectories seen so far by SGD. But even then the relation of $p(w|D_{n})$ with the distribution of SGD endpoints (with initialization and sampling providing the randomness) is obscure. This is the open question, mentioned in the introduction, of the relation between posterior concentration and SGD dynamics.
In this paper we take a more local perspective. The Bayesian antecedent hypothesis proposed in Chen et al. (2024b) claims that dynamical transitions (such as the ones observed in our experiments, where the optimization trajectory moves from plateau to plateau via sudden drops) are the “shadow” of Bayesian phase transitions in the following sense: each plateau corresponds to a region of parameter space around some local minima of the regret (a “phase”) and if there is a dynamical transition between two plateaus, there should be a Bayesian phase transition (the Bayesian antecedent) in which the posterior shifts concentration between the corresponding regions.
The existence of such a Bayesian antecedent does make a prediction: the LLC should rise during dynamical transitions. This was observed in Chen et al. (2024b) for supervised learning, and we see similar results here in the setting of reinforcement learning. In our experiments we can see that the estimate of the LLC coincides with a naïve assessment of policy complexity and phase transitions correspond to transitions from simpler but less optimal policies to policies that are more complex but more optimal. In the Bayesian setting singular learning theory, and the LLC in particular, provides a rigorous formalization of the notion of complexity of learning for deep RL models. Our empirical results are evidence that this notion of complexity is also applicable in the context of stochastic optimization.

### v2

@@ 5.2 The Generalized Free Energy Formula
where $U$ is a sufficiently small open neighborhood of $w_{0}$. This formula describes the concentration of the posterior in a region $U$ in terms of two invariants: the empirical regret $G_{n}(w_{0})$ (lower regret means lower $F_{n}(U)$, hence higher $Z_{n}(U)$ and thus posterior concentration) and the learning coefficient $\lambda$ (lower $\lambda$ means a simpler parametrized policy, and higher posterior concentration). Note that the prior $\phi(w)$ contributes to the free energy, and thus posterior concentration, only through constant order terms. These terms can certainly play a significant role for low $n$, but we do not consider this explicitly here.
5.3 Stagewise Development
The central insight for Bayesian deep RL that we draw from the free energy formula (14) is that learning consists of an evolving tradeoff between regret and complexity. This is the translation to RL of a perspective due to Watanabe (2009, §7.6; see also , ).
@@ 5.3 Stagewise Development
Note that since the relationship between the free energy and the posterior is exponential, these switches in the sign of the free energy as a function of $n$ can translate to rapid changes in posterior concentration. Following Watanabe (2018) and Chen et al. (2023), we refer to these sudden shifts in the concentration of the posterior as Bayesian phase transitions.
This suggests a picture of the Bayesian learning process as a series of phase transitions, where the posterior jumps between neighborhoods of local minima $w_{i}$ of $G$, according to a changing tradeoff between $G$ and $\lambda$ in the free energy formula. The most typical transition decreases $G$ and increases $\lambda$ (the posterior is “willing,” at some dataset size, to pay the free-energetic cost of higher complexity in exchange for lower regret).
5.4 Predictions for Deep Reinforcement Learning
Thus far we have given a formal description of the dynamics of the generalized Bayesian learning process in RL as a sequence of Bayesian phase transitions in which the generalized posterior $p(w|D_{n})\propto\exp(-nG_{n}(w))\phi(w)$ shifts its concentration between local minima of the population regret $G$. Here, $n$ denotes the number of trajectories “seen” by the learner and incorporated into the posterior update.
In SGD-based deep RL, the training process draws for each gradient step batches of $B$ trajectories. In a policy gradient algorithm like REINFORCE, the gradient estimator is essentially $\nabla_{w}G_{n}(w)$ with $n=B$. After $S$ steps of gradient descent, information from $B\cdot S$ trajectories have influenced the current parameter.
To attempt to line these two learning processes up as much as possible, take the sequence of behavioral parameters to be $w_{1},\ldots,w_{1},\ldots,w_{S},\ldots,w_{S}$ where $w_{i}$ is the $i$th step of gradient descent and each parameter is repeated $B$ times (cf., Section 3.2). Then, take the $\tau_{i}$ to be the contents of mini-batches sampled during SGD. $D_{n}$ is then the set of trajectories seen so far by SGD.
Note that the precise relationship between $p(w|D_{n})$ and, e.g., the distribution of SGD endpoints (with initialization and sampling providing the randomness) is yet to be formally characterized (see Section 7.3). However, we can take the theoretical phenomenon of Bayesian phase transitions and derive qualitative predictions about the dynamics of SGD-based RL, which can then be empirically tested:
  1. 1.
    We predict the existence of dynamical transitions in SGD-based deep RL, such that the optimization trajectory moves from plateau to plateau via sudden drops in the regret.
Stagewise development has been observed in deep RL before see, e.g., Clift et al., 2020; McGrath et al., 2022; also in neuroscience, e.g., Liebana et al., 2025. Our theory leads us to a more specific prediction:
  1. 2.
    The LLC should rise during these dynamical transitions, corresponding to an intuitive increase in policy complexity.
In summary, we predict that SGD-based deep RL should proceed stagewise from simple, high-regret policies to complex, low-regret policies, with policy complexity measured by the LLC.
These predictions parallel the situation in supervised learning. In particular, Chen et al. (2023) propose the Bayesian antecedent hypothesis, which claims that dynamical transitions are the “shadow” of Bayesian phase transitions in the following sense: each plateau corresponds to a region of parameter space around some local minima of the regret (a “phase”) and if there is a dynamical transition between two plateaus, there should be a corresponding Bayesian phase transition (the Bayesian antecedent) in which the posterior shifts concentration between the corresponding regions.
Evidence for dynamical transitions in which the LLC increases have been presented by Chen et al. (2023) in the case of small auto-encoders, and by Hoogland et al. (2025) and Carroll et al. (2025) for transformers trained on language data or synthetic in-context linear regression data.
@@ 6.1 Summary
During training we see stagewise development in policy space: the model transitions between plateaus in which the regret is roughly constant, the policy remains in a fixed region with qualitatively consistent behavior and the transitions between plateaus are relatively rapid. We refer to periods where the regret is approximately constant as phases and transitions between them as phase transitions (and one might compare this phenomenon to the phase transition seen in a toy model of supervised learning in Chen et al., 2023). For example, we will discuss some of the following common phases in policy space below.
@@
    We see gradual decrease in the LLC estimate within phases 2 and 3 – the policy, and therefore the regret, does not change significantly but the representation of the policy within parameter space does change. The LLC of phase 1 exhibits a consistent value of $29.54\pm 6.35$ during training across independent seeds.
@@ 6.2.2 Model
The architecture of the model directly follows that used for Abdel Sadek et al. (2025), which itself is a modified version of the Large architecture described in Espeholt et al. (2018): A convolutional network with 15 convolutional layers, but the LSTM that normally follows is replaced by a simple feedforward network that takes the output from the convolutional block. Full details of the architecture are given in Espeholt et al. (2018).
@@
Given the simplicity of the environment and given that there are only $121\times 120=14520$ states we can analytically compute the optimal reward $R_{\mathrm{max}}$ and the exact reward $R(w)$ of a given current policy $\pi_{w}$. This allows us to analytically compute the regret, $G(w)=R_{\mathrm{max}}-R(w)$ during various stages of training, as well as visualize the policy for each possible location of the agent given a goal location.
@@ 6.3 Details of LLC Estimation
We define an estimator for the local learning coefficient (see Section 5.1) using the preconditioned stochastic gradient Langevin dynamics (pSGLD) algorithm (Li et al., 2015) to estimate sampling from the Gibbs posterior. This combines RMSNorm-style adaptive step sizes with SGLD (Welling and Teh, 2011). Rather than imposing the sharp cutoff associated to the restriction to a neighborhood we use a Gaussian prior centered at $w^{*}$ with variance $\sigma^{2}$. The Gibbs posterior then takes the form
@@
The hyperparameters are the sample size $n$, the inverse temperature $\beta$, which controls the contribution of the regret, and the variance $\sigma^{2}$, which controls proximity to $w^{*}$}. For more details on these hyperparameters, see Watanabe (2013); Lau et al. (2025); Hoogland et al. (2025). In loc. cit. the inverse of the variance $\sigma^{2}$ is referred to as the localization strength. Details of the hyperparameters used for LLC estimation can be found in Appendix A.
@@ 6.4 Phase Transitions
A training run with $\alpha=0.6,\gamma=0.98$ is shown in Figure 1. We plot the regret, the difference between the value of the current policy, and that of an optimal policy, as a function of the number of environment steps seen during training. Three chosen points during training are illustrated with depictions of the learned policy in each phase.
@@ 6.4.1 Automatic Phase Detection
The value of $\delta$ is chosen to be smaller than the normalized distance between phase 1 and phase 2b in order to detect the transition between those phases. The precise choice of $\delta=0.15$ is not critical for the location of the phase transitions, as there is a range over which we could choose $\delta$ with only minor changes to the corresponding locations of phase transitions (see Appendix H).
@@ 6.4.2 Distribution of Phase Transitions
The training trajectory – and therefore the time steps at which phase transitions occur – are random variables, in this section we discuss their empirical distribution. Specifically we will present statistics in twelve cases: where the model was trained with $\alpha=0.4,0.5,0.6$ and $0.7$ and with $\gamma=0.97,0.98$ and $0.99$. In each hyperparameter case we trained models with these hyperparameters and $\sim 25$ independent random seeds. In Figure 3 we exhibit the distribution of the training steps at which phase transitions occur as a function of $\gamma$ for each of the four sampled $\alpha$-values. The models transition into phase 1 consistently early in training, the transition from phase 1 to phase 2b depends on $\gamma$ as we will discuss below.
@@
    The number of training steps before transitioning into $\pi^{1}$ is very low, and unrelated to the choice of $\alpha$$\gamma$ presented.
@@
    The higher the discount rate, the longer it takes for the model to converge to $\pi^{2b}$, if at all. For $\alpha,\gamma=0.6,0.99$ and $0.7,0.99$, the policy region $\pi^{2b}$ is never observed, but is skipped entirely, transitioning from $\pi^{1}\to\pi^{3}$ directly without visiting $\pi^{2b}$ first. For $\alpha,\gamma=0.7,0.98$, during training the policy would sometimes, but not always, skip $\pi^{2b}$ and transition directly from $\pi^{1}\to\pi^{3}$.
We can also try to predict the phase transitions based on the hyperparameters with which it was trained, see Appendix B.
@@ 6.4.3 LLC Estimates and Phases
Figure 1 shows the LLC estimate and regret across training for a representative model trained with $\alpha=0.6$ and $\gamma=0.98$, exhibiting two phase transitions. We present all the estimated LLC curves in Appendix F, Figure 7. We see rapid increases in the LLC estimate at each phase transition, and the increasingly complex phases are associated with increasingly large LLC estimates. We also note a gradual decrease of the LLC estimates in the second halves of phases 2 and 3 as stochastic gradient descent converges to simpler representations of the corresponding local minimum.
@@
Next we discuss the distribution of the LLC estimates within each phase. We performed LLC estimation for 25 independent models trained with $\alpha=0.6$ and $\gamma=0.98$ – those hyperparameters being selected to maximize the proportion of runs exhibiting two phase transitions. We estimated the LLC at six equally-logarithmically-spaced checkpoints in each of the three phases. Figure 4 shows the distribution of the LLC estimate at these points during training. Using the four internal points within each phase we see mean LLC estimates of 29.54 in phase one (with sample variance $\hat{\sigma}=6.35$), 125.53 in phase two ($\hat{\sigma}=46.47$) and 328.31 in phase three ($\hat{\sigma}=84.67$).
@@
Figure 4: Distribution of average LLC estimates across phases. The LLC has been estimated on distribution with $\alpha=0.6$ and $\gamma=0.98$.
@@ 7 Discussion
In this section, we discuss several implications of our results. In Section 7.1, we discuss some implication of stagewise learning dynamics for the important problem of AI alignment. In Section 7.2, we survey related work on model complexity and RL. In Section 7.3, we discuss open questions and new directions raised now that we have extended SLT to RL.